# On the Absolute-Value Integral of a Brownian Motion with Drift: Exact and Asymptotic Formulae

### v2

@@ 1 Introduction
With these fundamental properties, the process $X$ is a natural substitute for the closely related square integral functional $\big{(}\int^{t}_{0}\tilde{W}^{2}_{s}{\rm d}s\big{)}_{t\geq 0}$, corresponding to the $\mathrm{L}^{2}$-norm, whose distribution has been thoroughly studied in [40, Xia, 2020] (see also [31, Tolmatz, 2002] for the case of a standard Brownian bridge), and the exponential functional $\big{(}\int^{t}_{0}e^{\tilde{W}_{s}}{\rm d}s\big{)}_{t\geq 0}$, which has a long strand of literature (see, e.g., [22, Lyasoff, 2016] for a detailed analysis with the inclusion of drift). Specifically, $X$ finds applications in reliability engineering as a model for the random wear of structural materials [35, Van Noortwijk, 2009] and, in financial engineering, when establishing stochastic volatility based on the technique of stochastic time change ([4, Carr et al., 2003]) or when describing stochastic intensities in a reduced-form credit risk model ([9, Gourieroux et al., 2006]). It is worth noting that in the last aspect, the square integral has recently been adopted by [6, Chen and Xia, 2023+], which has ensured explicit evaluation of the cumulative intensity under measure changes – a very useful property that affine jump-diffusion models do not have; in connection with this, the drift component is crucial in generating sizable leverage effects. In comparison, the process $X$ tends to behave less variably than its square counterpart and is more suitable for describing calmer random phenomena for the same purposes. This is our main motivation to study (1.2).
This being said, the process $X$ also bears useful connections with various tools for statistical inference. For instance, the standard integral $\int^{1}_{0}|W_{s}|{\rm d}s$ is heavily tied to the so-called “empirical process” that appears in goodness-of-fit tests after applying the Khmaladze martingale transformation ([18, Khmaladze, 1993]), a technique designed to establish distributional independency when multiple parametric hypotheses are in place. In particular, the distribution of this integral is the (model-free) limit distribution of the corresponding test statistic based on the $\mathrm{L}^{1}$-norm, which is more robust to significant outliers than the commonly adopted statistics based on the $\mathrm{L}^{\infty}$-norm or the $\mathrm{L}^{2}$-norm; see, e.g., [27, Roberts et al., 2023] for illustrations of the latter two. The distribution function will also be explicitly given as a side result of our analysis. In the same direction, the general integral $X_{1}$ can also be identified as the (parameterizable) limit distribution of the conditional log-likelihood function when independent observations have Laplace distributions with normal prior means. We note that likelihood functions of this type sit at the heart of hierarchical matrix factorization for the exponential family; see, e.g., [36, Wang and Carvalho, 2023] for a recent development.
@@ 2 Laplace transform
×$\quad\;\times\bigg{(}\frac{\mathrm{\Gamma}(i)}{3^{i/3}\mathrm{\Gamma}(i/3)}+\frac{(-1)^{i}(-\alpha^{\prime}_{k-j})^{i+1}}{3^{1/3}}\bigg{(}\frac{1}{3^{1/3}\mathrm{\Gamma}(2/3)(i+1)}{\;{}_{1}\mathrm{F}_{2}}\bigg{(}\frac{i+1}{3};\frac{2}{3},\frac{i+4}{3}\bigg{|}\frac{(\alpha^{\prime}_{k-j})^{3}}{9}\bigg{)}$
@@
$\bar{f}_{X_{t}}(u)$ $=\int_{\mathds{R}}e^{bz-1/2b^{2}t}\sum^{\infty}_{k=1}\frac{v^{1/3}e^{(v^{2}/2)^{1/3}t\alpha^{\prime}_{k}}\mathrm{Ai}(\alpha^{\prime}_{k}+(2v)^{1/3}|z|)}{2^{2/3}(-\alpha^{\prime}_{k})\mathrm{Ai}(\alpha^{\prime}_{k})}{\rm d}z$
$=\sum^{\infty}_{k=1}\frac{v^{1/3}e^{(v^{2}/2)^{1/3}t\alpha^{\prime}_{k}-1/2b^{2}t}}{2^{2/3}(-\alpha^{\prime}_{k})\mathrm{Ai}(\alpha^{\prime}_{k})}\int_{\mathds{R}}e^{bz}\mathrm{Ai}(\alpha^{\prime}_{k}+(2v)^{1/3}|z|){\rm d}z.$ (2.5)
Since $\mathrm{Ai}$ is an entire function, by the substitutions ±$\pm(2v)^{1/3}z\mapsto z$ the last integral can be reformatted into
@@
$\bar{f}_{X_{t}}(u)=\sum^{\infty}_{k=1}\sum^{\infty}_{j=0}\frac{(b(2v)^{-1/3})^{2j}e^{(v^{2}/2)^{1/3}t\alpha^{\prime}_{k}-1/2b^{2}t}\mathcal{I}_{k,j}}{(-\alpha^{\prime}_{k})\mathrm{Ai}(\alpha^{\prime}_{k})(2j)!},\quad v=\sigma u,$
@@
An alternative way to obtain the formula (1) is by applying the celebrated Feynman–Kac formula and solving a partial differential equation that governs the function +×$\mathds{R}_{+}\times\mathds{R}\ni(t,z)\mapsto\varphi(t,z):=\mathbb{E}e^{-u\int^{t}_{0}|\mu s+\sigma(W_{s}+z)|{\rm d}s}$, with $u$ assumed to be positive-valued. More specifically, this equation reads
$\frac{\partial}{\partial t}\varphi(t,z)=\frac{1}{2}(\sigma u)^{2}\frac{\partial^{2}}{\partial z^{2}}\varphi(t,z)+\mu u\frac{\partial}{\partial z}\varphi(t,z)-|z|\varphi(t,z),$
@@ 3 Distribution functions
×$\quad\times\sum^{2j}_{i=0}\binom{2j}{i}(-\alpha^{\prime}_{k-j})^{2j-i}\bigg{(}\frac{\mathrm{\Gamma}(i)}{3^{i/3}\mathrm{\Gamma}(i/3)}+\frac{(-1)^{i}(-\alpha^{\prime}_{k-j})^{i+1}}{3^{1/3}}$
×$\qquad\times\bigg{(}\frac{1}{3^{1/3}\mathrm{\Gamma}(2/3)(i+1)}{\;{}_{1}\mathrm{F}_{2}}\bigg{(}\frac{i+1}{3};\frac{2}{3},\frac{i+4}{3}\bigg{|}\frac{(\alpha^{\prime}_{k-j})^{3}}{9}\bigg{)}+\frac{(-\alpha^{\prime}_{k-j})}{\mathrm{\Gamma}(1/3)(i+2)}{\;{}_{1}\mathrm{F}_{2}}\bigg{(}\frac{i+2}{3};\frac{4}{3},\frac{i+5}{3}\bigg{|}\frac{(\alpha^{\prime}_{k-j})^{3}}{9}\bigg{)}\bigg{)}$
×$\quad\times\mathrm{G}^{0,3}_{3,2}\bigg{(}\begin{array}[]{ccccc}(j+1)/3&(j+2)/3&j/3+1&&\\ &&&1/2&1\end{array}\bigg{|}\frac{27x^{2}}{2\sigma^{2}t^{3}(-\alpha^{\prime}_{k-j})^{3}}\bigg{)},\quad x>0,$ (3.4)
@@
×$\quad\times\mathrm{G}^{0,3}_{3,2}\bigg{(}\begin{array}[]{ccccc}(j+1)/3&(j+2)/3&j/3+1&&\\ &&&1/2&1\end{array}\bigg{|}\frac{27x^{2}}{2\sigma^{2}t^{3}(-\alpha^{\prime}_{k-j})^{3}}\bigg{)},\quad x>0,$ (3.9)
@@
×$\qquad\times\bigg{(}\frac{\mathrm{\Gamma}(i)}{3^{i/3}\mathrm{\Gamma}(i/3)}+\frac{(-1)^{i}(-\alpha^{\prime}_{k-j})^{i+1}}{3^{1/3}}\bigg{(}\frac{1}{3^{1/3}\mathrm{\Gamma}(2/3)(i+1)}{\;{}_{1}\mathrm{F}_{2}}\bigg{(}\frac{i+1}{3};\frac{2}{3},\frac{i+4}{3}\bigg{|}\frac{(\alpha^{\prime}_{k-j})^{3}}{9}\bigg{)}$
@@
×$\quad\times\mathrm{G}^{0,3}_{3,2}\bigg{(}\begin{array}[]{ccccc}j/3+5/6&j/3+7/6&j/3+3/2&&\\ &&&1/2&1\end{array}\bigg{|}\frac{27x^{2}}{2\sigma^{2}t^{3}(-\alpha^{\prime}_{k-j})^{3}}\bigg{)},\quad x>0.$ (3.14)
@@
$\quad\;+\frac{(-\alpha^{\prime}_{k})}{2\mathrm{\Gamma}(1/3)}{\;{}_{1}\mathrm{F}_{2}}\bigg{(}\frac{2}{3};\frac{4}{3},\frac{5}{3}\bigg{|}\frac{(\alpha^{\prime}_{k})^{3}}{9}\bigg{)}\bigg{)}\mathrm{G}^{0,3}_{3,2}\bigg{(}\begin{array}[]{ccccc}5/6&7/6&3/2&&\\ &&&1/2&1\end{array}\bigg{|}\frac{27x^{2}}{2\sigma^{2}t^{3}(-\alpha^{\prime}_{k-j})^{3}}\bigg{)},\quad x>0.$ (3.17)
@@ 4 Moments
×$\qquad\times\sum^{i}_{l=0}\beta^{l}\mathbf{B}_{i,l}\bigg{(}\frac{\mathrm{\Gamma}(5/3)}{\mathrm{\Gamma}(5/3-m)}y^{1-3m/2}\bigg{|}m\in\mathds{N}\cap[1,i-l+1]\bigg{)},$
@@
where the sum runs over all such combinations $\{\iota_{m}:m\in\mathds{N}\cap[1,i-l+1]\}$ that $\sum^{i+l-1}_{m=1}\iota_{m}=l$ and $\sum^{i+l-1}_{m=1}m\iota_{m}=i$. The same holds for $g^{(j)}_{-,k}(y^{3/2})$ except that $\beta$ is replaced by $-\beta$.
@@
×$\qquad\times\sum^{i}_{l=0}\beta^{l}\mathds{1}_{\text{even}}(l)\mathbf{B}_{i,l}\bigg{(}\frac{\mathrm{\Gamma}(5/3)}{\mathrm{\Gamma}(5/3-m)}y^{1-3m/2}\bigg{|}m\in\mathds{N}\cap[1,i-l+1]\bigg{)}$
@@
×$\qquad\times\frac{\mathrm{\Gamma}(1/2-k)}{\mathrm{\Gamma}(1/2-k-(l-i))}y^{-3/4-3(k+(l-i))/2}.$
@@
×$\quad\times\sum^{l}_{i=0}\binom{j}{i+j-l}\mathbf{B}_{j+i-l,j-l}\bigg{(}\frac{\mathrm{\Gamma}(5/3)}{\mathrm{\Gamma}(5/3-m)}y^{1-3m/2}\bigg{|}m\in\mathds{N}\cap[1,i+1]\bigg{)}\frac{\mathrm{\Gamma}(1/2-k)}{\mathrm{\Gamma}(1/2-k-(l-i))}y^{-3/4-3(k+(l-i))/2}$
@@
×$\quad\times\sum^{l}_{i=0}\binom{j+l}{j+i}\mathbf{B}_{j+i,j}\bigg{(}\frac{\mathrm{\Gamma}(5/3)}{\mathrm{\Gamma}(5/3-m)}y^{1-3m/2}\bigg{|}m\in\mathds{N}\cap[1,i+1]\bigg{)}\frac{\mathrm{\Gamma}(1/2-n+l)}{\mathrm{\Gamma}(1/2-n+i)}y^{-3/4-3(n-i)/2}.$ (4.11)
Thus, if we divide (4) by the product of $-2^{1/3}v^{4/3}$ and $\mathrm{Ai}^{\prime}(ay)$ as in (4.8), we obtain an asymptotic expansion of $\phi(u,y)$ in $y$ involving only nonnegative integer powers of $v=\sigma u$. More specifically, by applying Cauchy’s product rule and arranging terms we have
@@
×$\quad\times\sum^{l}_{i=0}\binom{j+l}{j+i}\mathbf{B}_{j+i,j}\bigg{(}\frac{\mathrm{\Gamma}(5/3)}{\mathrm{\Gamma}(5/3-m)}y^{1-3m/2}\bigg{|}m\in\mathds{N}\cap[1,i+1]\bigg{)}$
×$\qquad\times\frac{\mathrm{\Gamma}(1/2-k+l)}{\mathrm{\Gamma}(1/2-k+i)}y^{-1-3(k-i)/2}.$
@@
×$\quad\;\times\mathbf{B}_{2j+i,2j}\bigg{(}\frac{\mathrm{\Gamma}(5/3)}{\mathrm{\Gamma}(5/3-m)}y^{1-3m/2}\bigg{|}m\in\mathds{N}\cap[1,i+1]\bigg{)}y^{-1-3(n-i)/2}$ (4.13)
@@
×$\quad\times\sum^{\infty}_{j=0}\bigg{(}\frac{b^{2}}{2}\bigg{)}^{j}\frac{(2j+l)!}{(2j)!}P_{i}(j)y^{-j-1-3n/2}.$ (4.16)
@@
×$\quad\;\times\sum^{\infty}_{j=0}\bigg{(}\frac{\mu^{2}t}{2\sigma^{2}}\bigg{)}^{j}\frac{(2j+l)!}{(2j)!\mathrm{\Gamma}(j+3n/2+1)}P_{i}(j),$ (4.17)
@@
On another look, $\;{}_{2}\mathrm{F}_{2}$ in (4) only has integer or half-integer parameters, and so is always reducible to sums of the exponential and error functions. This means that, when fully simplified, all orders of moments should involve $\mathrm{erf}$ as the only non-elementary function. [^5] Further inspections indicate that even-order moments can be expressed free of non-elementary functions ($\mathrm{erf}$), which we prefer not to expound in this context. Achieving such a level of simplicity is completely a matter of collecting numerous terms of combined powers of the parameters $\mu$, $\sigma$, and $t$, which task becomes extremely tedious as soon as $n$ exceeds 2 but fortunately can be assigned to computer algebra. In particular, for $n\in\{0,1,2,3,4\}$, ${}^{\circledR}$ ([38, Wolfram Research, Inc., 2023]) calculates that ($\mu\neq 0$)
@@
[…]
@@
(iii) The skewness of $X_{t}$ goes to its infimum of 0 in the limit as $|\mu|\rightarrow\infty$, but there exists $\mu^{\ast}>0$ (depending on $\sigma$ and $t$) such that the excess kurtosis attains its global minimum [^7] Approximately $-0.2621350996$. at $\mu=\pm\mu^{\ast}$.

### v4

@@ 1 Introduction
With these fundamental properties, the process $X$ is a natural substitute for the closely related square integral functional $\big(\int^{t}_{0}\tilde{W}^{2}_{s}{\rm d}s\big)_{t\geq 0}$, corresponding to the $\mathrm{L}^{2}$-norm, whose distribution has been thoroughly studied in [40, Xia, 2020] (see also [31, Tolmatz, 2002] for the case of a standard Brownian bridge), and the exponential functional $\big(\int^{t}_{0}e^{\tilde{W}_{s}}{\rm d}s\big)_{t\geq 0}$, which has a long strand of literature (see, e.g., [22, Lyasoff, 2016] for a detailed analysis with the inclusion of drift). Specifically, $X$ finds applications in reliability engineering as a model for the random wear of structural materials [35, Van Noortwijk, 2009] and, in financial engineering, when establishing stochastic volatility based on the technique of stochastic time change ([4, Carr et al., 2003]) or when describing stochastic intensities in a reduced-form credit risk model ([9, Gourieroux et al., 2006]). It is worth noting that in the last aspect, the square integral has recently been adopted by [6, Chen et al., 2025], which has ensured explicit evaluation of the cumulative intensity under measure changes – a very useful property that affine jump-diffusion models do not have; in connection with this, the drift component is crucial in generating sizable leverage effects. In comparison, the process $X$ tends to behave less variably than its square counterpart and is more suitable for describing calmer random phenomena for the same purposes. This is our main motivation to study (1.2).
This being said, the process $X$ also bears useful connections with various tools for statistical inference. For instance, the standard integral $\int^{1}_{0}|W_{s}|{\rm d}s$ is heavily tied to the so-called “empirical process” that appears in goodness-of-fit tests after applying the Khmaladze martingale transformation ([18, Khmaladze, 1993]), a technique designed to establish distributional independency when multiple parametric hypotheses are in place. In particular, the distribution of this integral is the (model-free) limit distribution of the corresponding test statistic based on the $\mathrm{L}^{1}$-norm, which is more robust to significant outliers than the commonly adopted statistics based on the $\mathrm{L}^{\infty}$-norm or the $\mathrm{L}^{2}$-norm; see, e.g., [27, Roberts et al., 2025] for illustrations of the latter two. The distribution function will also be explicitly given as a side result of our analysis. In the same direction, the general integral $X_{1}$ can also be identified as the (parameterizable) limit distribution of the conditional log-likelihood function when independent observations have Laplace distributions with normal prior means. We note that likelihood functions of this type sit at the heart of hierarchical matrix factorization for the exponential family; see, e.g., [36, Wang and Carvalho, 2023] for a recent development.
@@ 2 Laplace transform
$\quad\;\times\bigg(\frac{\mathrm{\Gamma}(i)}{3^{i/3}\mathrm{\Gamma}(i/3)}+\frac{(-1)^{i}(-\alpha^{\prime}_{k-j})^{i+1}}{3^{1/3}}\bigg(\frac{1}{3^{1/3}\mathrm{\Gamma}(2/3)(i+1)}{\;{}_{1}\mathrm{F}_{2}}\bigg(\frac{i+1}{3};\frac{2}{3},\frac{i+4}{3}\bigg|\frac{(\alpha^{\prime}_{k-j})^{3}}{9}\bigg)$
@@
$\bar{f}_{X_{t}}(u)$ $=\int_{\mathds{R}}e^{bz-b^{2}t/2}\sum^{\infty}_{k=1}\frac{v^{1/3}e^{(v^{2}/2)^{1/3}t\alpha^{\prime}_{k}}\mathrm{Ai}(\alpha^{\prime}_{k}+(2v)^{1/3}|z|)}{2^{2/3}(-\alpha^{\prime}_{k})\mathrm{Ai}(\alpha^{\prime}_{k})}{\rm d}z$
$=\sum^{\infty}_{k=1}\frac{v^{1/3}e^{(v^{2}/2)^{1/3}t\alpha^{\prime}_{k}-b^{2}t/2}}{2^{2/3}(-\alpha^{\prime}_{k})\mathrm{Ai}(\alpha^{\prime}_{k})}\int_{\mathds{R}}e^{bz}\mathrm{Ai}(\alpha^{\prime}_{k}+(2v)^{1/3}|z|){\rm d}z.$ (2.5)
Since $\mathrm{Ai}$ is an entire function, by the substitutions $\pm(2v)^{1/3}z\mapsto z$ the last integral can be reformatted into
@@
$\bar{f}_{X_{t}}(u)=\sum^{\infty}_{k=1}\sum^{\infty}_{j=0}\frac{(b(2v)^{-1/3})^{2j}e^{(v^{2}/2)^{1/3}t\alpha^{\prime}_{k}-b^{2}t/2}\mathcal{I}_{k,j}}{(-\alpha^{\prime}_{k})\mathrm{Ai}(\alpha^{\prime}_{k})(2j)!},\quad v=\sigma u,$
@@
An alternative way to obtain the formula (1) is by applying the celebrated Feynman–Kac formula and solving a partial differential equation that governs the function $\mathds{R}_{+}\times\mathds{R}\ni(t,z)\mapsto\varphi(t,z):=\mathbb{E}e^{-u\int^{t}_{0}|\mu s+\sigma(W_{s}+z)|{\rm d}s}$. More specifically, this equation reads
$\frac{\partial}{\partial t}\varphi(t,z)=\frac{1}{2}\sigma^{2}\frac{\partial^{2}}{\partial z^{2}}\varphi(t,z)+\mu\frac{\partial}{\partial z}\varphi(t,z)-u|z|\varphi(t,z),$
@@ 3 Distribution functions
$\quad\times\sum^{2j}_{i=0}\binom{2j}{i}(-\alpha^{\prime}_{k-j})^{2j-i}\bigg(\frac{\mathrm{\Gamma}(i)}{3^{i/3}\mathrm{\Gamma}(i/3)}+\frac{(-1)^{i}(-\alpha^{\prime}_{k-j})^{i+1}}{3^{1/3}}$
$\qquad\times\bigg(\frac{1}{3^{1/3}\mathrm{\Gamma}(2/3)(i+1)}{\;{}_{1}\mathrm{F}_{2}}\bigg(\frac{i+1}{3};\frac{2}{3},\frac{i+4}{3}\bigg|\frac{(\alpha^{\prime}_{k-j})^{3}}{9}\bigg)+\frac{(-\alpha^{\prime}_{k-j})}{\mathrm{\Gamma}(1/3)(i+2)}{\;{}_{1}\mathrm{F}_{2}}\bigg(\frac{i+2}{3};\frac{4}{3},\frac{i+5}{3}\bigg|\frac{(\alpha^{\prime}_{k-j})^{3}}{9}\bigg)\bigg)$
$\quad\times\mathrm{G}^{0,3}_{3,2}\bigg(\begin{array}[]{ccccc}(j+1)/3&(j+2)/3&j/3+1&&\\ &&&1/2&1\end{array}\bigg|\frac{27x^{2}}{2\sigma^{2}t^{3}(-\alpha^{\prime}_{k-j})^{3}}\bigg),\quad x>0,$ (3.4)
@@
$\quad\times\mathrm{G}^{0,3}_{3,2}\bigg(\begin{array}[]{ccccc}(j+1)/3&(j+2)/3&j/3+1&&\\ &&&1/2&1\end{array}\bigg|\frac{27x^{2}}{2\sigma^{2}t^{3}(-\alpha^{\prime}_{k-j})^{3}}\bigg),\quad x>0,$ (3.9)
@@
$\qquad\times\bigg(\frac{\mathrm{\Gamma}(i)}{3^{i/3}\mathrm{\Gamma}(i/3)}+\frac{(-1)^{i}(-\alpha^{\prime}_{k-j})^{i+1}}{3^{1/3}}\bigg(\frac{1}{3^{1/3}\mathrm{\Gamma}(2/3)(i+1)}{\;{}_{1}\mathrm{F}_{2}}\bigg(\frac{i+1}{3};\frac{2}{3},\frac{i+4}{3}\bigg|\frac{(\alpha^{\prime}_{k-j})^{3}}{9}\bigg)$
@@
$\quad\times\mathrm{G}^{0,3}_{3,2}\bigg(\begin{array}[]{ccccc}j/3+5/6&j/3+7/6&j/3+3/2&&\\ &&&1/2&1\end{array}\bigg|\frac{27x^{2}}{2\sigma^{2}t^{3}(-\alpha^{\prime}_{k-j})^{3}}\bigg),\quad x>0.$ (3.14)
@@
$\quad\;+\frac{(-\alpha^{\prime}_{k})}{2\mathrm{\Gamma}(1/3)}{\;{}_{1}\mathrm{F}_{2}}\bigg(\frac{2}{3};\frac{4}{3},\frac{5}{3}\bigg|\frac{(\alpha^{\prime}_{k})^{3}}{9}\bigg)\bigg)\mathrm{G}^{0,3}_{3,2}\bigg(\begin{array}[]{ccccc}5/6&7/6&3/2&&\\ &&&1/2&1\end{array}\bigg|\frac{27x^{2}}{2\sigma^{2}t^{3}(-\alpha^{\prime}_{k})^{3}}\bigg),\quad x>0.$ (3.17)
@@ 4 Moments
$\qquad\times\sum^{i}_{l=0}\beta^{l}\mathbf{B}_{i,l}\bigg(\frac{\mathrm{\Gamma}(5/3)}{\mathrm{\Gamma}(5/3-m)}y^{1-3m/2}\bigg|m\in\mathds{N}\cap[1,i-l+1]\bigg),$
@@
where the sum runs over all such combinations $\{\iota_{m}:m\in\mathds{N}\cap[1,i-l+1]\}$ that $\sum^{i+l-1}_{m=1}\iota_{m}=l$ and $\sum^{i+l-1}_{m=1}m\iota_{m}=i$. The same holds for $g^{(j)}_{-,k}(y^{3/2})$ except that $\beta$ is replaced by $\beta$.
@@
$\qquad\times\sum^{i}_{l=0}\beta^{l}\mathds{1}_{\text{even}}(l)\mathbf{B}_{i,l}\bigg(\frac{\mathrm{\Gamma}(5/3)}{\mathrm{\Gamma}(5/3-m)}y^{1-3m/2}\bigg|m\in\mathds{N}\cap[1,i-l+1]\bigg)$
@@
$\qquad\times\frac{\mathrm{\Gamma}(1/2-k)}{\mathrm{\Gamma}(1/2-k-(l-i))}y^{-3/4-3(k+(l-i))/2}.$
@@
$\quad\times\sum^{l}_{i=0}\binom{j}{i+j-l}\mathbf{B}_{j+i-l,j-l}\bigg(\frac{\mathrm{\Gamma}(5/3)}{\mathrm{\Gamma}(5/3-m)}y^{1-3m/2}\bigg|m\in\mathds{N}\cap[1,i+1]\bigg)\frac{\mathrm{\Gamma}(1/2-k)}{\mathrm{\Gamma}(1/2-k-(l-i))}y^{-3/4-3(k+(l-i))/2}$
@@
$\quad\times\sum^{l}_{i=0}\binom{j+l}{j+i}\mathbf{B}_{j+i,j}\bigg(\frac{\mathrm{\Gamma}(5/3)}{\mathrm{\Gamma}(5/3-m)}y^{1-3m/2}\bigg|m\in\mathds{N}\cap[1,i+1]\bigg)\frac{\mathrm{\Gamma}(1/2-n+l)}{\mathrm{\Gamma}(1/2-n+i)}y^{-3/4-3(n-i)/2}.$ (4.11)
Thus, if we divide (4) by the product of $2^{1/3}v^{4/3}$ and $\mathrm{Ai}^{\prime}(ay)$ as in (4.8), we obtain an asymptotic expansion of $\phi(u,y)$ in $y$ involving only nonnegative integer powers of $v=\sigma u$. More specifically, by applying Cauchy’s product rule and arranging terms we have
@@
$\quad\times\sum^{l}_{i=0}\binom{j+l}{j+i}\mathbf{B}_{j+i,j}\bigg(\frac{\mathrm{\Gamma}(5/3)}{\mathrm{\Gamma}(5/3-m)}y^{1-3m/2}\bigg|m\in\mathds{N}\cap[1,i+1]\bigg)$
$\qquad\times\frac{\mathrm{\Gamma}(1/2-k+l)}{\mathrm{\Gamma}(1/2-k+i)}y^{-1-3(k-i)/2}.$
@@
$\quad\;\times\mathbf{B}_{2j+i,2j}\bigg(\frac{\mathrm{\Gamma}(5/3)}{\mathrm{\Gamma}(5/3-m)}y^{1-3m/2}\bigg|m\in\mathds{N}\cap[1,i+1]\bigg)y^{-1-3(n-i)/2}$ (4.13)
@@
$\quad\times\sum^{\infty}_{j=0}\bigg(\frac{b^{2}}{2}\bigg)^{j}\frac{(2j+l)!}{(2j)!}P_{i}(j)y^{-j-1-3n/2}.$ (4.16)
@@
$\quad\;\times\sum^{\infty}_{j=0}\bigg(\frac{\mu^{2}t}{2\sigma^{2}}\bigg)^{j}\frac{(2j+l)!}{(2j)!\mathrm{\Gamma}(j+3n/2+1)}P_{i}(j),$ (4.17)
@@
On another look, $_{2}\mathrm{F}_{2}$ in (4) only has integer or half-integer parameters and so is always reducible to sums of the exponential and error functions. This means that, when fully simplified, all orders of moments should involve $\mathrm{erf}$ as the only non-elementary function. [^5] Further inspections indicate that even-order moments can be expressed free of non-elementary functions ($\mathrm{erf}$), which we prefer not to expound in this context. Achieving such a level of simplicity is completely a matter of collecting numerous terms of combined powers of the parameters $\mu$, $\sigma$, and $t$, which task becomes extremely tedious as soon as $n$ exceeds 2 but fortunately can be assigned to computer algebra. In particular, for $n\in\{0,1,2,3,4\}$, Mathematica® ([38, Wolfram Research, Inc., 2023]) calculates that ($\mu\neq 0$)
@@
$\max_{\mu\in\mathds{R}}\mathrm{Skew}X_{t}=\frac{8(4480-1257\pi)}{35\sqrt{(27\pi-64)^{3}}}\quad\text{and}\quad\max_{\mu\in\mathds{R}}\mathrm{EKurt}X_{t}=\frac{3(858112\pi-33453\pi^{2}-2293760)}{280(27\pi-64)^{2}}.$
@@
(iii) The skewness of $X_{t}$ goes to its infimum of 0 in the limit as $|\mu|\rightarrow\infty$, but there exists $\mu^{\ast}>0$ (depending on $\sigma$ and $t$) such that the excess kurtosis attains its global minimum [^7] Approximately 0.2621350996-0.2621350996. at $\mu=\pm\mu^{\ast}$.